The lifetime X of an alkaline battery is exponentially distributed with lambda = 0.05 per hour. What are the mean and standard deviation of the battery’s lifetime


1
Expert's answer
2022-05-11T17:44:08-0400

If a random variable "X" is exponentially distributed with rate "\u03bb," then


"E(X)=\dfrac{1}{\lambda}, Var(X)=\sigma^2=(\dfrac{1}{\lambda})^2"

Given "\lambda=0.05." Then


"E(X)=\dfrac{1}{0.05}=20, \sigma=\sqrt{\sigma^2}=\dfrac{1}{0.05}=20"